### v1

# Towards Optimal Distributed Edge Coloring with Small Palettes

Abstract
We design a deterministic distributed $\mathcal{O}(\log n)$-round reduction from the $(2\Delta-2)$-edge coloring problem to the much easier $(2\Delta-1)$-edge coloring problem. This is almost optimal, as the $(2\Delta-2)$-edge coloring problem admits an $\Omega(\log_{\Delta}n)$ lower bound. Further, we also obtain an optimal $\mathcal{O}(\log_{\Delta}n)$-round reduction, albeit to the harder maximal independent set (MIS) problem. In our approach we find almost edge-degree-choosable clusters of small diameter. The edges inside these clusters will be colored last by our algorithm. More precisely, we show that if we allow each cluster of uncolored edges to change the color of just two edges in its $1$-hop neighborhood, then it can extend any partial $(2\Delta-2)$-edge coloring. In order to perform these color changes in parallel we need to assign independent edges to each cluster. Therefore we construct an edge set that satisfies two key properties: (1) Each cluster needs to receive exclusive access to two edges in its $1$-hop neighborhood. (2) The set of all such edges needs to be a matching.
Computing a greedy maximal matching satisfies (2), but not necessarily (1). Hence, we modify the maximal matching by letting each cluster send proposals to matching edges in its $2$-hop neighborhood. If the number of proposals that each cluster sends exceeds the number of proposal each matching edge receives we can efficiently assign two matching edges to each cluster. Each cluster can then move its assigned edges into its $1$-hop neighborhood, satisfying (1), while maintaining the matching property (2). This is the only “non-greedy” part of our reduction and it contributes the bulk of the runtime. Finally, each cluster can change the colors of its assigned edges in parallel such that it can extend the coloring to the uncolored edges inside the cluster.
The current state-of-the-art for $(2\Delta-1)$-edge coloring actually comes from an MIS algorithm by [Ghaffari & Grunau, FOCS’24], which runs in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds. With our new reduction, this round complexity now carries over to the $(2\Delta-2)$-edge coloring problem as well. Alternatively, one can also plug in the $(\mathrm{poly}\log\Delta+\mathcal{O}(\log^{\ast}n))$-round $(2\Delta-1)$-edge coloring algorithm from [Balliu, Brandt, Kuhn & Olivetti, PODC’22], which yields an optimal runtime of $\mathcal{O}(\log n)$ rounds for $\Delta\leq\mathrm{poly}\log n$. Previously, the fastest deterministic algorithm using less than $12\Delta-$ 1 colors for general graphs by [Brandt, Maus, Narayanan, Schager & Uitto, SODA’25] ran in $\widetilde{\mathcal{O}}(\log^{3}n)$ rounds. In addition, we also obtain a $\mathcal{O}(\log\log n)$-round randomized reduction of $(2\Delta-2)$-edge coloring to $(2\Delta-1)$-edge coloring. This improves upon the (very recent) best randomized algorithm using less than $12\Delta-$ 1 colors from [Bourreau, Brandt & Nolin, STOC’25] by reducing the round complexity from $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ down to $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$.
keywords:
distributed graph algorithms, edge coloring, $\mathsf{LOCAL}$ model
1 Introduction
In this work we present distributed algorithms for the well-known edge coloring problem. Edge coloring is considered to be one of the four classical problems in the area of local distributed algorithms, together with vertex coloring, maximal independent set (MIS) and maximal matching. A $k$-edge coloring of a graph $G=(V,E)$ is a function $\varphi:E\to\{1,\dots,k\}$ such that no two edges sharing an endpoint receive the same color. The parameter $k$ is called the palette size. Another important parameter for graph coloring problems is the maximum degree $\Delta$. In the centralized setting an edge coloring that uses at least $12\Delta-$ 1 colors can be computed by a trivial greedy algorithm: Simply process all the edges in an arbitrary order and pick any available color at every step. Since any edge is adjacent to most $22\Delta-$ 2 other edges, at least one color will always be available. Similar greedy algorithms also work for the classical $(\Delta+1)$-vertex coloring, MIS and maximal matching problems. Therefore, we will henceforth refer to problems of this type as greedy problems. In contrast, for any palette size $k<2\Delta-1$ there are partial $k$-edge colorings that cannot be extended to the entire graph. In particular, in the online setting, where the input is given as a stream of edges or vertices and the algorithm has to make irrevocable decisions based on partial information on the graph, there is no algorithm that can edge color any graph with less than $12\Delta-$ 1 colors due to a seminal result by Bar-Noy, Motwani and Naor [2]. Therefore this so-called “greedy color threshold” $k=2\Delta-1$ presents a natural and important barrier in the development of distributed algorithms for this problem.
How much more difficult are non-greedy problems than greedy problems?
Going below $12\Delta-$ 1 colors requires completely new techniques, as the algorithm needs to ensure that the yet uncolored edges do not lose too many available colors. This additional difficulty is also demonstrated by the fact that coloring the edges of a graph with less than $12\Delta-$ 1 colors admits a $\Omega(\log_{\Delta}n)$-round deterministic lower bound [9] in the $\mathsf{LOCAL}$ model, while there are deterministic $\mathsf{LOCAL}$ algorithms that compute a $(2\Delta-1)$-edge coloring in $\mathcal{O}(\log^{\ast}n)$ rounds [3, 35], at least in graphs of constant maximum degree. It is known for a long time that every graph can be edge-colored using just $\Delta+1$ colors [43], however until very recently there was no way to compute such a coloring efficiently in the distributed setting.
Model of computation
Our model of computation is the classical message-passing model of distributed computation introduced by Linial in the 1990’s [32, 33]. Today it is known under the apt name $\mathsf{LOCAL}$, coined by Peleg in [40]. In the $\mathsf{LOCAL}$ model the communication network is abstracted as an undirected graph. The vertices are processors with unlimited local computation power and globally unique identifiers. They can communicate with their neighbors in synchronous rounds. The message size is assumed to be unlimited and the complexity of an algorithm is defined as the number of synchronous round until every node outputs (a local part of) the solution. The related $\mathsf{CONGEST}$ model additionally imposes a maximum message size of $\mathcal{O}(\log n)$ bits per message.
Greedy distributed edge coloring
One of the first results for edge coloring in the distributed setting can be derived from Luby’s celebrated MIS algorithm [34]. Via known reductions from edge coloring to vertex coloring to MIS this leads to a randomized algorithm for $(2\Delta-1)$-edge coloring running in $\mathcal{O}(\log n)$ rounds of the $\mathsf{LOCAL}$ model. Further, early works by Panconesi & Srinivasan [36, 39] show that using network decompositions, one can find a $(2\Delta-1)$-edge coloring in $2^{\mathcal{O}(\sqrt{\log n})}$ rounds deterministically. For graphs with constant maximum degree $\Delta$, one can even solve the problem in $\mathcal{O}(\log^{\ast}n)$ rounds [35], matching the famous $\Omega(\log^{\ast}n)$ lower bound by Linial [32, 33]. For general graphs, however, the hunt for the first polylogarithmic-time algorithm that uses just $12\Delta-$ 1 colors was an important and long-standing open question in the field until it was finally positively answered by Fischer, Ghaffari & Kuhn in 2017 [15]. The key technical ingredient of their seminal $\mathcal{O}(\log^{8}\Delta\cdot\log n)$-round algorithm is a deterministic algorithm for the hypergraph maximal matching problem. Since then, the field has picked up the pace with a lot of new algorithms being developed in the last decade, e.g. [1, 19, 20, 21, 25, 29, 30]. The state-of-the-art $(2\Delta-1)$-edge coloring for general graphs now comes out of an MIS algorithm developed in [20]. It runs in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds deterministic and $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds randomized time.
Breaking the greedy threshold
While most of the algorithms that break the greedy color barrier were only developed in the last ten years, there are two remarkable results before the turn of the millennium. In a seminal paper from 1992, Panconesi & Srinivasan [37, 38] presented the first distributed algorithm for the $\Delta$-vertex coloring algorithm. Their algorithm is deterministic and runs in $2^{\mathcal{O}(\sqrt{\log n})}$ rounds of the $\mathsf{LOCAL}$ model. Since the edge coloring problem on a graph of maximum degree $\Delta$ is equivalent to coloring the vertices of the corresponding line graph of maximum degree $22\Delta-$ 2, this result also yields the first edge-coloring algorithm below the greedy color threshold. On the randomized side, in 1998, Dubhashi, Grable & Panconesi [12] presented an $(1+\varepsilon)\Delta$-edge coloring algorithm running in $\mathcal{O}(\log n)$ rounds, provided that $\Delta>(\log n)^{1+\gamma}$, using the Rödl nibble method. Subsequently, both the number of rounds, as well as the palette size has undergone significant further improvements, which we will discuss in detail in Section 1.3. To put our new results into context we will just briefly mention the relevant state-of-the art algorithms for $(2\Delta-2)$-edge coloring here. In a recent paper studying the hypergraph sinkless orientation problem [8], the authors present a deterministic $(3/2+\varepsilon)\Delta$-edge coloring algorithm running in $\widetilde{\mathcal{O}}(\varepsilon^{-2}\cdot\log^{2}\Delta\cdot\log n)$ rounds. This algorithm is still the fastest for general graphs, even if we allow the palette size to grow as large as $22\Delta-$ 2. On the randomized side, the fastest $(2\Delta-2)$-edge coloring algorithm works by computing a $\Delta$-vertex coloring in $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds [7].
1.1 Our contributions
[…]
Theorem 1.1.
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathcal{O}(\log_{\Delta}n)+T_{\mathrm{MIS}}(\Delta^{2}\cdot n,\mathrm{poly}(\Delta))$ rounds of the $\mathsf{LOCAL}$ model.
By plugging in the state-of-the-art MIS algorithm from [20] we get the following runtime.
Corollary 1.2.
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds of the $\mathsf{LOCAL}$ model.
Secondly, via replacing the expensive MIS by a cheaper ruling set we also get a slightly slower $\mathcal{O}(\log n)$-round reduction to the much easier $(2\Delta-1)$-edge coloring problem.
Theorem 1.3.
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathcal{O}(\log n)+T_{2\Delta-1}(n,\Delta)$ rounds of the $\mathsf{LOCAL}$ model.
Now we can also plug in the $(\mathrm{poly}\log\Delta+\mathcal{O}(\log^{\ast}n))$-round $(2\Delta-1)$-edge coloring algorithm from [1] to get an almost optimal runtime for $\Delta\leq\mathrm{poly}\log n$.
Corollary 1.4.
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathrm{poly}\log\Delta+\mathcal{O}(\log n)$ rounds of the $\mathsf{LOCAL}$ model.
Further, we show that our deterministic result can be modified to also yield a faster randomized reduction. Since our main technical contributions are already fast enough, this only requires us to replace the deterministic subroutines for ruling sets, maximal matchings, etc. by their (sometimes exponentially) faster randomized counterparts.
Theorem 1.5.
[…]
Again, plugging in the state-of-the-art randomized $(2\Delta-1)$-edge coloring algorithm from [20] yields the following explicit runtime.
Corollary 1.6.
There exists a randomized distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds of the $\mathsf{LOCAL}$ model.
1.2 Our technique in a nutshell
(a) This coloring can be extended.
(b) This coloring cannot be extended.
Figure 1: Extendability of $(2\Delta-2)$-colorings for $\Delta=3$.
The core observation forming the heart of our reduction is that partial $(2\Delta-2)$-edge colorings can almost always be extended. To give some intuition for this fact, consider the simplest possible case, where the set of uncolored edges in relation to a partial $(2\Delta-2)$-edge coloring $\varphi$ forms a star graph $S\subseteq G$. The only case, in which $\varphi$ cannot be extended to $S$ is if each edge in $S$ has the exact same set of $\Delta-1$ colors appearing among its adjacent edges, which we illustrate in Figure 1. An interesting application of Hall’s marriage theorem shows that this is indeed the only bad case. Next, we will show that this condition generalizes to trees as well. The core observation of this second step is that as long as the leaves in layer $k$ of the tree see at least $\Delta$ distinct colors among their incident edges, we can extend the coloring such that the leaves in layer $k-1$ see at least $\Delta$ distinct colors as well, see Figure 2. Hence, it is sufficient to have one layer of the tree that satisfies this colorful condition.
Lemma 1.7 (Simplified version of Lemma 3.4).
Consider a $(2\Delta-2)$-edge colored graph except for a complete $\Delta$-ary subtree $T\subseteq G$. The coloring can be extended to $T$, if the leaves of $T$ have at least $\Delta$ colors appearing on their incident edges combined.
In the remainder of this paper we will refer to the above condition as the colorful condition. In the full version of this lemma, we generalize the statement to any subtree of a $\Delta$-regular graph, which requires a little bit more care in how one formulates the condition of having enough colors available for the leaves. The main technical difficulty to overcome is that now the leaves of $T$ can all be at different distances from the root.
The reason why this condition is so powerful is because it does not require much intervention in order to fulfil it. For a $\Delta$-regular graph $G$, all we need to do is to change the color of one single edge adjacent to the cluster. Since each leaf in $T$ is already adjacent to $\Delta-1$ colored edges with necessarily distinct colors we only need to obtain one single additional color. However, there are still two big challenges ahead of us, if we want to make this approach work. Firstly, we need to ensure that the chosen edge actually has a useful free color to switch to. Secondly, we need to ensure that all clusters can perform this color switching in parallel without interfering with each other.
To overcome the first obstacle, we construct a subgraph $G^{\prime}\subseteq G$ of maximum degree at most $\Delta-1$ that contains all edges of $G$ except for a spanning tree of each cluster. This subgraph can then be $(2\Delta-3)$-edge coloring using a regular $(2\Delta-1)$-edge coloring algorithm. Hence, every edge in $E(G^{\prime})$ has at least the color $22\Delta-$ 2 available.
To address the second obstacle, we need that the set of edges selected by each cluster is independent, that is, it forms a matching in $G^{\prime}$. Also, for technical reasons we will assign two edges to each cluster instead of just one. Each cluster will only modify the color of one of those two edges, but it needs to ensure that the color of the second edge remains unchanged. As a good first step towards our goal we construct a maximal matching $M$ of $G^{\prime}$, which is still a greedy problem. This is not yet enough though, as the maximality of $M$ only guarantees that each vertex on the boundary of a given cluster sees at least one matching edge in its $2$-hop neighborhood. Hence, there might be a cluster entirely without adjacent matching edges.
Therefore our next step is to modify this matching to meet the above criterion. To solve this problem, we first let the clusters compete for the current edges in $M$. Each cluster $C$ wants to obtain exclusive access to two edges from $M$, which it can then shift along a shortest path towards $C$ until the edge is adjacent to $C$. Hence, each cluster selects a sufficiently large set of vertices on its boundary and lets each such vertex send a proposal to one of its closest matching edges. Note that each edge $e\in M$ may receive proposals from up to $2\Delta$ different clusters. Hence, it suggests itself to model this problem as a hypergraph $H$, where the vertices are the clusters in $G$. Each edge $e\in M$ defines a hyperedge by collecting all clusters that sent a proposals to $e$. What we need to solve now is the non-greedy assignment problem of hyperedges (matching edges) to vertices (clusters). Luckily, this problem has already been studied in the literature, where it is known under the name hypergraph sinkless orientation. An orientation of a hypergraph $H$ defines each hyperedge to be outgoing for exactly one vertex and ingoing for all others. A sinkless orientation on $H$ then requires that every vertex $v$ has at least one outgoing hyperedge, which we can then assign to $v$. As long as each cluster sends out significantly more proposals than any edge in $M$ receives, this problem can be solved in $\mathcal{O}(\log n)$ rounds deterministically and $\mathcal{O}(\log\log n)$ rounds randomized [8]. Finally, each cluster now has two edges, whose colors it can control independently of all other clusters. In Lemma 4.16 we formally show that this is indeed enough to fulfil the colorful condition in Lemma 3.4.
(a) The edges with different colors palettes available are dashed.
(b) The coloring can be extended while using at least three colors for the newly colored edges.
Figure 2: Extending the coloring to a tree.
1.3 Related work
While the majority of the distributed community has previously focused on developing algorithms using at least $12\Delta-$ 1 colors [1, 4, 13, 14, 15, 19, 25, 20, 29, 34, 35], there has been some notable progress below the greedy threshold. We already discussed some of the early works, which showed how to compute a $(2\Delta-2)$-edge coloring deterministically [37, 38] or a $(1+\varepsilon)\Delta$-edge coloring randomized [12]. On the lower bound side, the authors of [9] established that coloring the edges of a graph with at most $22\Delta-$ 2 colors takes at least $\Omega(\log_{\Delta}n)$ rounds deterministically and $\Omega(\log_{\Delta}\log n)$ rounds randomized. In Table 1 we attempt to give a concise overview over the algorithms using less than $12\Delta-$ 1 colors.
Randomized edge coloring via the distributed Lovász Local Lemma
After the turn of the millennium, the attention towards this problem dwindled down for a while, with the next significant progress appearing not until 2015 from Elkin, Pettie & Su [13]. They improve upon [12] by providing a randomized $(1+\varepsilon)\Delta$-edge coloring algorithm that reduces the problem to the distributed Lovász Local Lemma and runs in $\mathcal{O}(\log^{\ast}\Delta+\frac{\log n}{\Delta^{1-o(1)}})$ rounds. In both algorithms it is clear that $\varepsilon$ need not be constant, but it is not obvious how small it can be made as a function of $\Delta$. Optimizing for palette size, the authors of [9] give a randomized $(\Delta+\widetilde{\Theta}(\sqrt{\Delta}))$-edge coloring algorithm that runs in $\mathrm{poly}(\log n)$ rounds. This almost reaches the natural threshold $\Delta+\Theta(\sqrt{\Delta})$ for randomized coloring algorithms, which is the smallest palette size for which there still is a constant probability of being able to color any edge given a random feasible coloring of its neighborhood. Further improvements to variants of the distributed Lovász Local Lemma [28, 11] subsequently reduced the randomized complexity of $(1+\varepsilon)\Delta$-edge coloring further down to $\mathrm{poly}(\log\log n)$.
Deterministic edge coloring
In [26] the authors leverage maximum matchings in bipartite graphs to get a $\mathcal{O}(\Delta^{9}\cdot\mathrm{poly}(\log n))$-time deterministic algorithm for $(3\Delta/2)$-edge coloring. Their technique involves iteratively removing graphs of maximum degree $3$, which reduce the maximum degree of the original graph by at least $2$. By coloring each of those extracted graphs with a distinct set of three colors they end up with a $(3\Delta/2)$-edge coloring of the entire graph. This result has undergone a series of improvements since then. First, Harris [30] presented improved algorithms for hypergraph maximal matchings, culminating in a reduced runtime of $\widetilde{\mathcal{O}}(\Delta^{4}\cdot\log^{6}n)$ rounds. Next, via a novel algorithm for the hypergraph sinkless orientation problem [8] the complexity was pushed down even further to just $\mathcal{O}(\Delta^{2}\cdot\log n)$ rounds. Furthermore, by leveraging a splitting technique from [24, 23] this algorithm can also be turned into an $(3/2+\varepsilon)\Delta$-edge coloring algorithm running in $\widetilde{\mathcal{O}}(\varepsilon^{-2}\cdot\log^{2}\Delta\cdot\log n)$ rounds. In particular, this implies a $\mathcal{O}(\log^{3}n)$-round algorithm for $(2\Delta-2)$-edge coloring in general graphs, which is still the state-of-the-art even for this larger palette size. Another technique to obtain deterministic edge coloring algorithms that use fewer colors is the method of conditional expectation, which can be used to obtain a $(\Delta+\mathrm{poly}(\log n))$-edge coloring algorithm running in $\Delta+\mathrm{poly}(\log n)$ rounds [21].
$\Delta$-vertex coloring
Building on the seminal work in [37], the authors of [22] decreased the upper bound for the randomized complexity of the $\Delta$-coloring problem down to $\mathcal{O}(\log\Delta)+2^{\mathcal{O}(\sqrt{\log\log n})}$ after over two decades of no progress on this problem. Further, a very recent result by Bourreau, Brandt and Nolin [7] shows that $\Delta$-coloring can be solved in $\mathcal{O}(\log^{4}\Delta+\log^{2}\Delta\cdot\log n\log^{\ast}n)$ deterministic time and $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ time randomized. This also automatically yields corresponding $(2\Delta-2)$-edge coloring algorithms. While the time complexity of the deterministic algorithm does not yield an improvement over the $(3/2+\varepsilon)\Delta$-edge coloring algorithm from [8], the randomized version is still the current state-of-the-art for $(2\Delta-2)$-edge coloring. Finally we note that for $\Delta=\omega(\log^{21}n)$ the randomized complexity even collapses to $\mathcal{O}(\log^{\ast}n)$ due to [16].
Edge coloring via Vizing chains
Vizing’s theorem [43] states that every simple undirected graph of maximum degree $\Delta$ can be edge-colored using just $\Delta+1$ colors. Since $\Delta$ is an easy lower bound on the number of colors necessary, the edge-chromatic number of any graph falls into one of just two classes. In the last years there has been a considerable amount of work dedicated to distributed versions of Vizing’s theorem. In [42], the authors give a randomized $\mathrm{poly}(\Delta,\log n)$ algorithm for $(\Delta+2)$-edge coloring as well as a deterministic $\mathrm{poly}(\Delta,\log n)$ algorithm that uses $\Delta+\mathcal{O}(\log n/\log\log n)$ colors. In [6] the authors give a deterministic $(\Delta+1)$-edge coloring algorithm in $\widetilde{\mathcal{O}}(\mathrm{poly}(\Delta)\cdot\log^{5}n)$ rounds of the $\mathsf{LOCAL}$ model, as well as a $\mathcal{O}(\mathrm{poly}(\Delta)\cdot\log^{2}n)$-round randomized algorithm. Optimizing for palette size, there is a recent deterministic $(\Delta+1)$-edge coloring algorithm [5, 6] in $\widetilde{\mathcal{O}}(\Delta^{84}\cdot\log^{5}n)$ rounds. While the exponent in $\Delta$ in the runtime is certainly not optimized yet, there is a fundamental limit to this approach. The authors of [9] have shown that this approach requires recoloring subgraphs of diameter $\Omega(\Delta\cdot\log n)$ in the worst case.
Citation Year Runtime # Colors Random?
[37] 1992 $2^{\mathcal{O}(\sqrt{\log n})}$ $22\Delta-$ 2
[21] 2018 $\mathrm{poly}(\Delta,\log n)$ $\Delta+\mathrm{poly}(\log n)$
[26] 2018 $\mathcal{O}(\Delta^{9}\cdot\mathrm{poly}\log n)$ $23\Delta/$ 2
[29] 2019 $\widetilde{\mathcal{O}}(\Delta^{4}\cdot\log^{6}n)$ $23\Delta/$ 2
[42] 2019 $\mathrm{poly}(\Delta,\log n)$ $\Delta+\mathcal{O}(\log n/\log\log n)$
[5, 6] 2022 $\widetilde{\mathcal{O}}(\Delta^{84}\cdot\log^{5}n)$ $\Delta+1$
[8] 2025 $\mathcal{O}(\Delta^{2}\cdot\log n)$ $23\Delta/$ 2
[8] 2025 $\widetilde{\mathcal{O}}(\varepsilon^{-2}\cdot\log^{2}\Delta\cdot\log n)$ $(3/2+\varepsilon)\Delta$
[7] 2025 $\mathcal{O}(\log^{4}\Delta+\log^{2}\Delta\cdot\log n\log^{\ast}n)$ $22\Delta-$ 2
[12] 1998 $\mathcal{O}(\log n)$ $(1+\varepsilon)\Delta$dice6
[13] 2015 $\mathcal{O}(\log^{\ast}\Delta\cdot\lceil\frac{\log n}{\Delta^{1-o(1)}}\rceil)$ $(1+\varepsilon)\Delta$dice6
[…]
[9] 2018 $\mathrm{poly}(\log n)$ $\Delta+\widetilde{\Theta}\left(\sqrt{\Delta}\right)$dice6
[42] 2019 $\mathrm{poly}(\Delta,\log n)$ $\Delta+2$dice6
[28] 2022 $\mathrm{poly}(\log\log n)$ $(1+\varepsilon)\Delta$dice6
[5, 6] 2023 $\mathcal{O}(\mathrm{poly}(\Delta)\cdot\log^{2}n)$ $\Delta+1$dice6
[11] 2023 $\mathrm{poly}(\varepsilon^{-1}\cdot\log\log n)$ $(1+\varepsilon)\Delta$dice6
[…]
Table 1: Edge-coloring algorithms in the LOCAL model using fewer than $12\Delta-$ 1 colors.
1.4 Organization of the paper
In Section 3 we present the main technical condition under which partial $(2\Delta-2)$-edge colorings can be extended. We start Section 4 by presenting the simpler version of our new algorithm, which provides the reduction to MIS. Later on, in Section 4.5, we explain which subroutines can be replaced their relaxed counterparts with faster algorithms in order to achieve a reduction to $(2\Delta-1)$-edge coloring. Finally, in Section 5 we explain which subroutines need to be replaced in order to obtain the faster randomized reduction. We use the appendix to give a short overview over all the subroutines from the existing literature used throughout our algorithms to be used as a point of reference.
2 Preliminaries
We will use the notation $[k]:=\{1,\dots,k\}$. Let $G=(V,E)$ be an undirected graph. The neighborhood of a vertex $v\in V$ in $G$ is $N_{G}(v)=\{w\in V:vw\in E\}$ and the edge-neighborhood of $v$ is $N_{E}(v)=\{e\in E:\lvert e\cap\{v\}\rvert=1\}$. Further, for a subset of vertices $W\subseteq V$ we write $N_{E}(W)=\{e\in E:\lvert e\cap W\rvert=1\}$. The distance between two vertices $u,v\in V$ is defined as the number of edges on a shortest path connecting $u$ and $v$ in $G$ and is denoted by $\mathrm{dist}(u,v)$. We define the distance between an edge $e=uv\in E$ and a vertex $w\in V$ as
$\mathrm{dist}(e,w):=\min(\mathrm{dist}(u,w),\mathrm{dist}(v,w)).$
For any positive integer $k>0$ we define the power graph $G^{k}=(V,E^{k})$ by adding an edge for every pair of vertices $u,v\in V$ with $\mathrm{dist}(u,v)\leq k$ to $E^{k}$. For any positive integer $k>0$ and any vertex $v\in V$ we define the $r$-hop neighborhood of $v$ to be $N^{k}(v)=N_{G^{k}}(v)$. For a subset of edges $F\subseteq E$ we write $G-F$ to denote the graph obtained by removing all edges in $F$ from $G$. Further, for a subset $V^{\prime}\subseteq V$, we denote by $G[V^{\prime}]=(V^{\prime},E^{\prime})$ the subgraph of $G$ induced by $V^{\prime}$. The edge set $E^{\prime}$ is given by the subset of edges with both endpoints in $V$. Similarly, for a subset of edges $E^{\prime}\subseteq E$ we define the edge-induced subgraph $G[E^{\prime}]=(V^{\prime},E^{\prime})$, where $V^{\prime}$ is given by the subset of vertices that are incident to at least one edge in $E^{\prime}$. We define the degree $\mathrm{deg}(e)$ of an edge $e\in E$ as the number of edges sharing an endpoint with $e$. A hypergraph is a generalization of a graph in which each edge can contain more than two vertices. The rank of a hyperedge $e$ is the number of vertices contained in $e$. A proper $c$-edge coloring is a function $\varphi:E\to[c]$ such that for all adjacent edges $e,e^{\prime}\in E$ it holds that $\varphi(e)\neq\varphi(e^{\prime})$. In this paper we will work a lot with partial edge colorings that assign colors only to a subset of edges.
Definition 2.1.
Let $G=(V,E)$ be a graph, $F\subseteq E$ and $\varphi:E\setminus F\to[2\Delta-2]$ a partial edge coloring of $G$. Then we call $\varphi^{\star}$ an extension of $\varphi$ to $F$ if $\varphi^{\star}$ is a proper $(2\Delta-2)$-edge coloring of $G$ and $\varphi^{\star}(e)=\varphi(e)$ for all $e\in E\setminus F$.
We say a color $c$ is available for an edge $e\in F$, if no adjacent edge already has color $c$.
Definition 2.2 (edge-degree-choosable).
A subgraph $H\subseteq G$ is edge-degree-choosable, if for every assignment of lists =$\mathcal{L}=\{L(e):e\in E(H)\}$, such that $|L(e)|\geq\deg(e)$ for all $e\in E(H)$, there exists a proper edge-coloring of $G$ with colors from $\mathcal{L}$.
Note that any edge-degree choosable subgraph $H\subseteq G$ can be $(2\Delta-2)$-edge colored no matter how the edges outside of $G$ are colored.
Definition 2.3 ($(\alpha,\beta)$-ruling set).
For two integers $\alpha,\beta\geq 1$ an $(\alpha,\beta)$-ruling set for a subset $W\subseteq V$ is a set of nodes $S\subseteq V$ such that the distance between any two nodes in $S$ is at least $\alpha$ and any node $w\in W$ has a distance of at most $\beta$ to the closest node in $S$. If $V=W$, we call $S$ an $(\alpha,\beta)$-ruling set.
Ruling sets generalize maximal independent sets, which are $(2,1)$-ruling sets. In our algorithm we will use ruling sets as a tool in order to partition the vertex sets into clusters with some specific properties. Most importantly, we will need that each cluster $C$ there is a central vertex $r\in C$ such that the whole $k$-hop neighborhood around $r$ is contained in $C$, for some parameter $k>0$. This ensures that each cluster has sufficiently many vertices, which will become important when we try to assign two exclusive matching edges to each cluster.
Definition 2.4 ($(\alpha,\beta)$-clustering).
Let $G=(V,E)$ be a graph and $\mathcal{R}\subseteq V$ a subset of vertices. Then, we call a partition $\bigcup_{r\in\mathcal{R}}C(r)=V$ of the vertex set into pairwise disjoint subsets a $(\alpha,\beta)$-clustering with respect to $\mathcal{R}$ if $N^{\alpha}(r)\subseteq C(r)$ and $\mathrm{diam}(C(r))\leq\beta$ for all $r\in\mathcal{R}$.
Remark 2.5.
Computing a $(2,\beta)$-ruling set on the power graph $G^{k}$ and letting each vertex join the cluster of one of its closest ruling set nodes leads to a $(\lfloor\frac{k}{2}\rfloor,k\cdot\beta)$-clustering.
In the third step of our algorithm, our goal is to assign two exclusive edges to each cluster, such that the set of all such edges forms a matching in $G$. Starting with a maximal matching, we construct an auxiliary hypergraph $H$ that lets us modify the matching in order to satisfy this property. The vertices of the hypergraph are the clusters in $G$ and we want to make sure that each cluster gets exclusive access to two hyperedges. This can be seen as an instance of the hypergraph sinkless orientation problem, which is a natural generalization of the fundamental sinkless orientation problem.
Definition 2.6 (Hypergraph sinkless orientation (HSO)).
Let $H=(V,E)$ be a hypergraph. The objective of the hypergraph sinkless orientation problem is to orient the hyperedges of $H$ such that every vertex $v\in V$ has at least one outgoing hyperedge. We define an oriented hyperedge to be outgoing for exactly one of its incident vertices and incoming for all others.
If the minimum degree $\delta=\min_{v\in V}\mathrm{deg}(v)$ and the maximum rank $r=\max_{e\in E}\mathrm{rank}(e)$ of a hypergraph $H=(V,E)$ satisfy $\delta>r$, then this problem can be solved efficiently using an algorithm from [8].
3 Under which condition can we extend partial $(2\Delta-2)$-colorings?
Our algorithm starts by computing a $(\alpha,\beta)$-clustering $\mathcal{C}$ with respect to some independent set $\mathcal{R}$ (the parameters $\alpha$ and $\beta$ will be fixed when they become relevant) by letting each vertex join the cluster of one of its closest nodes in $\mathcal{R}$. For each cluster $C(r)\in\mathcal{C}$, the central vertex $r\in\mathcal{R}$ saves the BFS-tree $T(r)$ that naturally arises from this clustering. Let $T=\bigcup_{r\in\mathcal{R}}T(r)$ be the forest obtained by joining all those BFS-trees. The edges in $T$ will be colored last by our algorithm. The remaining subgraph $G-E(T)$ has maximum degree at most $\Delta-1$ (as every vertex is part of exactly one such BFS-tree). Hence, we can invoke any regular $(2\Delta-1)$-edge coloring algorithm to get a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$. Therefore, the goal of this section is to establish under which conditions we can extend $\varphi$ to a proper $(2\Delta-2)$-edge coloring of the entire graph $G$. As the first step towards this goal we consider the case where the subgraph induced by the uncolored edges forms a simple star graph. In this case, the only forbidden configuration is if all leaves of the star graph are incident to the same set of $\Delta-1$ colors, which can be seen in Figure 1(b).
Lemma 3.1 (Reaching for the stars).
Let $G=(V,E)$ be a graph and let $\varphi$ be a partial $(2\Delta-2)$-edge coloring of $G$ that only leaves the edges of a star graph $T=(V_{T},E_{T})\subseteq G$ uncolored. If $\lvert\varphi(N_{E}(V_{T}))\rvert\geq\Delta$, then there is an extension $\varphi^{\star}$ of $\varphi$ to $E_{T}$.
To prove this lemma we invoke the following classical combinatorial result.
Theorem 3.2 (Hall’s theorem [27]).
Let $G=(V\,\dot{\cup}\,U,E)$ be a bipartite graph. For any subset $S\subseteq U$, let $N(S)$ denote the neighborhood of $S$ in $G$. Then $G$ has a $U$-saturating matching, if and only if $|N(S)|\geq|S|$ for all $S\subseteq U$.
Proof 3.3 (Proof of Lemma 3.1).
We construct the auxiliary bipartite graph $B=(E_{T}\cup[2\Delta-2],F)$ as follows: For every edge $e\in E_{T}$ and every color $c\in[2\Delta-2]$ we add the edge $(e,c)$ to $F$ if and only if $c$ is available for $e$ under $\varphi$. Next we observe that any $E_{T}$-saturating matching in $B$ corresponds to an extension of $\varphi$ to $E_{T}$. By Theorem 3.2 such a matching exists if and only if for any subset $S\subseteq E_{T}$ it holds that $\lvert N_{B}(S)\rvert\geq\lvert S\rvert$. Since any edge in $T$ has at least $\Delta-1$ colors available and $\lvert E_{T}\rvert\leq\Delta$, the only set left to check is $S=E_{T}$. For contradiction sake, assume that $\lvert N_{B}(E_{T})\rvert=\Delta-1$ and let $r\in V_{T}$ be the center of the star. This implies that for every $e=rv\in E_{T}$ all colors in $C:=[2\Delta-2]\setminus N_{B}(E_{T})$ are already occupied. Since $\lvert C\rvert=\Delta-1\geq\lvert N_{E}(v)\cap N_{E}(V_{T})\rvert$, this is only possible if $\varphi(N_{E}(v)\cap N_{E}(V_{T}))=C$ for all $e\in E_{T}$. But $\lvert\varphi(N_{E}(V_{T}))\rvert\geq\Delta$, a contradiction!
As the next step we show that this condition generalizes to trees as well. More precisely, we will show that if $\varphi$ assigns at least $\Delta$ distinct colors to the set of edges which are adjacent to at least one vertex at distance $k$ from the root, then there is always an extension of $\varphi$ that assigns at least $\Delta$ distinct colors to the edges adjacent to any vertex at distance $k-1$ from the root.
Lemma 3.4 (Colorful leaves make any tree happy).
Let $G=(V,E)$ be a $\Delta$-regular graph, $T=(V_{T},E_{T})\subseteq G$ a tree, $r\in V_{T}$ and $\varphi:E\setminus E_{T}\to[2\Delta-2]$ a partial edge coloring of $G$. Let
$V_{k}:=\{e\in V_{T}:\mathrm{dist}_{T}(v,r)=k\}\qquad\text{and}\qquad E_{k}=\{e\in E_{T}:\mathrm{dist}_{T}(e,r)=k\}.$
If there is a $k\in\mathbb{N}$ such that $|\varphi(N_{E}(V_{k}))|\geq\Delta$, then there exists an extension $\varphi^{\star}$ of $\varphi$ to $E_{T}$.
Proof 3.5.
Let $\ell\in\mathbb{N}$ such that $|\varphi(N_{E}(V_{\ell}))|\geq\Delta$. First we observe that for every $k>0$, each edge $e\in E_{k}$ has exactly one neighboring edge in $E_{k-1}$. Therefore, we can always find at least one available color for every such edge, as long as the set $E_{k-1}$ remains uncolored. Using this observation we can assign arbitrary available colors to all edges in $\bigcup_{k\geq\ell}E_{k}$ by processing these edge sets in sequence, starting with the set with highest index and going through each of these sets in an arbitrary order. However, after we finish coloring $E_{\ell}$ we need to start being more deliberate with our color choices. For any $k<\ell$ we need to ensure that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ holds after we colored all edges of $E_{k}$. This is due to the fact that we need $\lvert\varphi(N_{E}(V_{1}))\rvert\geq\Delta$ in order to apply Lemma 3.1, which guarantees that we can complete the coloring on the final layer of edges $E_{0}$ connected to the root. Now we show that given $\lvert\varphi(N_{E}(V_{k+1}))\rvert\geq\Delta$ we can color the edges in $E_{k}$ in such a way that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ for any $1\leq k<\ell$. For each $e\in E_{k}$ let $N^{+}(e):=N_{E}(e)\cap N_{E}(V_{k+1})$. Since $\lvert N^{+}(e)\rvert=\Delta-1$ we can always find two edges $e,e^{\prime}\in E_{k}$ such that $\varphi(N^{+}(e))\neq\varphi(N^{+}(e^{\prime}))$. We first assign colors to all edges in $E_{k}$ that are not adjacent to $e$ or $e^{\prime}$. Since we assume $\Delta$ to be sufficiently large, there always exists a vertex $v\in V_{k}$ that is not incident to either $e$ or $e^{\prime}$. Thus, after coloring all $\Delta-1$ edges in $N_{E}(v)\setminus E_{k-1}$ incident at $v$ we already have $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta-1$. If $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$, we are already done. Otherwise, if $\lvert\varphi(N_{E}(V_{k}))\rvert=\Delta-1$, then the set $\varphi(N_{E}(V_{k})\cup(N^{+}(e)\cap N^{+}(e^{\prime})))$ contains at most $32\Delta-$ 3 colors. Therefore, either $e$ or $e^{\prime}$ must have a color available in $[2\Delta-2]\setminus\varphi(N_{E}(V_{k}))$. We pick that color for $e$ or $e^{\prime}$, which ensures that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$. The remaining edges in $E_{k}$ can then be assigned arbitrary available colors again. Finally, to color the last layer $E_{0}$ containing the edges incident to the root, we can apply Lemma 3.1, since the number of distinct adjacent colors is at least $\Delta$.
Now in order to fulfil this condition, our algorithm assigns two exclusive edges to each cluster. Each cluster may then decide independently, whether it wants to keep the current color or change it to the (yet unused) color $22\Delta-$ 2 for each assigned edge. Hence, the set of all those edges needs to be independent. Therefore we initially compute a maximal matching on $G-E_{T}$ and modify it in a way such that each cluster receives at least two edges. We model this adjustment of $M$ as an instance of hypergraph sinkless orientation, where each vertex proposes to a suitable matching edge in its $2$-hop neighborhood. The core requirement for this problem is that the minimum degree on the vertex side must exceed the maximum rank on the hyperedge side. In other words, each cluster must send more proposals to distinct edges than the number of proposals that any edge receives. The details of this procedure will be explained in Section 4.3. Therefore we need the clusters to either expand quickly or not need such edges by virtue of being already edge-degree-choosable. For the latter case, we observe as soon as a connected component contains one cycle of even length or a vertex of degree less than $\Delta$, it is already edge-degree choosable.
Lemma 3.6.
Let $G=(V,E)$ and $\varphi:E\to[2\Delta-2]\cup\{\bot\}$ be a partial $(2\Delta-2)$-edge coloring of $G$. Assume that the subgraph $H\subseteq G$ induced by the uncolored edges is connected and contains either a cycle of even length or a vertex of degree less than $\Delta$. Then there exists an extension $\varphi^{\star}$ of $\varphi$, which is a proper $(2\Delta-2)$-edge coloring of the entire graph $G$.
Proof 3.7.
Let $C$ be a cycle of even length and assume that all edges in $N(C)$ have been colored already. Our task is now to find a proper edge-coloring of $C$, where every edge has a list of available colors of size two. If all edges have the same two colors available, then we can color the cycle alternately with these two colors. Otherwise, there exist two adjacent edges $e_{1},e_{2}$ with different sets of available colors. Color $e_{1}$ with a color that is not available for $e_{2}$. Then we traverse the cycle starting at $e_{1}$ and going in the opposite direction of $e_{2}$ and let each edge pick one of its available colors. Since $e_{1}$ picked a color that is not available at $e_{2}$, there is also a color remaining for $e_{2}$ at the end. Since $H$ is connected, every edge lies on a path towards an even cycle. Fix such an even cycle $C$ and define
$E_{k}=\{e\in E(H):\mathrm{dist}(e,C)=k\}.$
Since every edge in $E_{k}$ has at least one neighbor in $E_{k-1}$ for $k>1$, we can greedily assign colors to these edges, starting from the highest layer. Finally, every edge in $C$ has a list of two available colors and can therefore be properly colored as well.
If a cluster does not contain a cycle of even length, we need to make sure that it can satisfy the colorful condition from Lemma 3.4. For that we need to make sure that the cluster contains a sufficient number of leaves. We apply a similar argument to the one used in the analysis of the $\Delta$-vertex coloring algorithm in [22] to show that a connected component that does not contain an even cycle must be expanding exponentially.
Lemma 3.8 (Cluster expansion lemma).
Let $G=(V,E)$ be a $\Delta$-regular graph, $k>0,v\in V$ and $C\subseteq G$ be a connected subgraph of $G$ such that $N^{k}(v)\subseteq C$. Further assume that $C$ does not contain an cycle of even length. Then, any BFS-tree $T$ of $C$ rooted at $v$ contains at least $(\Delta-2)^{k}$ leaves.
Proof 3.9.
For two nodes $u,u^{\prime}\in V(C)$ let $P(u,u^{\prime})$ denote the unique path in $T$ between $u$ and $u^{\prime}$. We claim that each node $u\in V(C)$ is incident to at most one edge that is not part of $E(T)$ and thus the degree of all internal nodes of $T$ is at least $\Delta-1$. First we observe that any edge in $E(C)\setminus E(T)$ can only connect vertices which are at most one level apart in $T$. Any additional edge in $E(C)\setminus E(T)$ from a vertex $v_{\ell}$ in layer $\ell$ to a vertex $v_{\ell+1}$ in layer $\ell+1$ in $T$ immediately induces an even cycle. To wit, let $u$ be the least common ancestor of $v_{\ell}$ and $v_{\ell+1}$ and observe that the cycle $P(u,v_{\ell})\cup\{v_{\ell},v_{\ell+1}\}\cup P(u,v_{\ell+1})$ has length $2\cdot\lvert P(u,v_{\ell+1})\rvert$. Next, let $u\in V(C)$ be a vertex in layer $\ell$ with two neighbors $w,w^{\prime}$ in the same layer $\ell$. Then, $P(v,w)\cup\{w,u\}\cup\{u,w^{\prime}\}\cup P(v,w^{\prime})$ is a cycle of length $2(\ell+1)$. Hence each vertex $u\in V(C)$ is adjacent to at most one edge from $E(C)\setminus E(T)$. Since $C$ contains the complete $k$-hop neighborhood of $v$ this implies that there are at least $(\Delta-2)^{k}$ vertices in layer $k$ of $T$. Every such vertex either has a leaf-descendant in $T$ or is a leaf itself.
4 Deterministic $(2\Delta-2)$-edge coloring algorithm
We first present the simpler version of our algorithm, which reduces the $(2\Delta-2)$-edge coloring problem deterministically to MIS in $\mathcal{O}(\log_{\Delta}n)$ rounds of the $\mathsf{LOCAL}$ model.
4.1 High level overview of our reduction to MIS
Algorithm 1 Reduction from $(2\Delta-2)$-edge coloring to MIS (high level overview)
1:Input: a graph $G=(V,E)$
2:Output: a $(2\Delta-2)$-edge coloring $\varphi$ of $G$.
3:
4:Phase 1: Partition vertices into clusters
5:Compute an MIS $\mathcal{I}$ on the power graph $G^{8}$triangleright Theorem B.10
6:$\mathcal{C},E_{T}\leftarrow$ Cluster$(G,\mathcal{I})$triangleright Algorithm 2
7:Compute a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$triangleright Lemma 4.3
8:Phase 2: Assign two exclusive edges to each cluster
9:Compute a maximal matching $M$ of $G-E_{T}$triangleright Lemma 4.5
10:Set up auxiliary hypergraph $H=(\mathcal{C},M)$triangleright Definition 4.9
11:Assign two matching edges to each cluster via HSO on $H$triangleright Theorem B.8
12:Phase 3: Switch colors in order to complete the coloring
13:for each cluster in parallel do
14:     Change the colors of the assigned edges to fulfil the colorful condition $\triangleright$ Lemma 4.16
15:end for
16:Extend $\varphi^{\prime}$ to a $(2\Delta-2)$-edge coloring $\varphi^{\star}$ of $G$triangleright Lemma 3.4
Phase 1: Partition vertices into clusters (Section 4.2).
We start by computing a maximal independent set $\mathcal{I}$ on the power graph $G^{8}$. Next we partition the vertex set into disjoint clusters. Each vertex joins the cluster of one of its closest MIS-nodes, breaking ties arbitrarily, which leads to a $(4,8)$-clustering $\mathcal{C}$. In addition, we store the BFS-tree $(V,E_{T})$ which arises naturally from this clustering. Computing the MIS on $G^{8}$ ensures that the individual clusters all contain sufficiently many vertices, which will become important later. Next, we remove the union of all BFS-trees from $G$ to obtain a subgraph $G-E_{T}\subseteq G$ of maximum degree at most $\Delta-1$. Hence, it suffices to invoke any $(2\Delta-1)$-edge coloring on $G-E_{T}$ in order to get a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$.
Phase 2: Assign two exclusive edges to each cluster (Section 4.3).
Finally we need to modify $\varphi$ such that the colorful condition from Lemma 3.4 is satisfied for every cluster. We will show that this can be achieved by giving each cluster control over the colors of two edges adjacent to the cluster. However, in order for the clusters to switch the colors of their respective assigned edges, the union of all those edge sets needs to be independent. Therefore we initially compute a maximal matching $M$ of $G-E_{T}$ to use as a starting point. Note that the maximality of the matching does not guarantee that any fixed cluster receives any adjacent matching edges, but only that there are sufficiently many matching edges in the $2$-hop neighborhood around that cluster. Hence, we then let each cluster propose to matching edges in its $2$-hop neighborhood. This process naturally defines a hypergraph, where each matching edges collects the clusters which sent a proposal to it as vertices in a hyperedge. Then, the problem of assigning matching edges to clusters can be modeled as an instance of hypergraph sinkless orientation. In order to be able to solve HSO efficiently we need to ensure that the number of votes each matching edge receives is dominated by the number of votes each cluster casts in total. Here we require that each cluster contains enough vertices.
Phase 3: Switch colors in order to complete the coloring (Section 4.4).
Finally, we show that each cluster can flip the colors of their assigned matching edges in parallel in order to fulfil the colorful condition from Lemma 3.4. Thus, this new coloring can now be extended to the remaining uncolored edges by iteratively coloring the layers of each BFS-tree in parallel. Since each cluster has diameter of at most $8$ this last step can be executed in constant time.
See 1.1
4.2 Phase 1: Partition vertices into clusters
Algorithm 2 Cluster$(G,\mathcal{I})$
1:Input: a graph $G=(V,E)$ and an independent set $\mathcal{I}\subseteq V$
2:Output: a $(4,8)$-clustering $\mathcal{C}=\{C^{-1}(r):r\in\mathcal{I}\}$ and a forest $T=(V,E_{T})$
3:
4:$E_{T}\leftarrow\emptyset$
5:for each $v\in V$ in parallel do
6:     $C(v)\leftarrow\bot$
7:     if $v\in\mathcal{I}$ then
8:         $C(v)\leftarrow v$
9:     end if
10:     while $C(v)=\bot$ do
11:         receive messages from neighbors
12:         if $\exists\,u\in N(v):C(u)\neq\bot$ then
13:              $C(v)\leftarrow C(u)$
14:              $E_{T}\leftarrow E_{T}\cup\{uv\}$triangleright break ties arbitrarily
15:              send message $C(v)$ to neighbors
16:         end if
17:     end while
18:end for
Lemma 4.1 (Clustering algorithm).
The set $\mathcal{C}=\{C^{-1}(r):r\in\mathcal{I}\}$ computed by Algorithm 2 is a $(4,8)$-clustering with respect to $\mathcal{I}$.
Proof 4.2.
First we show that $N^{4}(r)\subseteq C^{-1}(r)$ for all $r\in\mathcal{I}$. Let $r\in\mathcal{I},v\in N^{4}(r)$ and assume for contradiction sake that $C(v)=r^{\prime}\neq r$. Then, $\mathrm{dist}(v,r^{\prime})\leq\mathrm{dist}(v,r)\leq 4$ and therefore $\mathrm{dist}(r,r^{\prime})\leq 8$, contradicting the fact that $r$ and $r^{\prime}$ are independent in $G^{8}$.
For the second property assume that there exist $r\in\mathcal{I}$ and $v\in C^{-1}(r)$ such that $\mathrm{dist}(v,r)>8$. This implies that $\mathrm{dist}(v,r^{\prime})\geq\mathrm{dist}(v,r)>8$ for all $r^{\prime}\in\mathcal{I}$, because otherwise $v$ would have joined the cluster of $r^{\prime}$ instead of $r$. But then $v$ would be independent of all $r^{\prime}\in\mathcal{I}$ on $G^{8}$, contradicting the maximality of $\mathcal{I}$!
For each $r\in\mathcal{I}$, let $T(r)=T[C^{-1}(r)]$ denote the BFS-tree of the cluster $C^{-1}(r)$ rooted at $r$.
Lemma 4.3.
The subgraph $G-E_{T}$ has maximum degree of at most $\Delta-1$ and can be $(2\Delta-3)$-edge colored in $T_{\mathrm{MIS}}(\Delta^{2}\cdot n,4\Delta)$ rounds of the $\mathsf{LOCAL}$ model.
Proof 4.4.
We observe that the edge set $E_{T}$ returned by Algorithm 2 satisfies $E_{T}=\bigcup_{r\in\mathcal{I}}E(T(r))$. For every vertex $v\in V$ there is an $r\in\mathcal{R}$ such that $v\in C^{-1}(r)$. Since $T(r)$ is a spanning tree of $C^{-1}(r)$, we get that $v$ is incident to at least one edge in $E(T(r))$. Thus, the maximum degree of $G-E_{T}$ is at most $\Delta-1$. Using the standard reduction from vertex coloring to MIS we can compute a $(2\Delta-3)$-edge coloring of $G-E_{T}$ in $T_{\mathrm{MIS}}(\Delta^{2}\cdot n,4\Delta)$ rounds.
Lemma 4.5.
We can compute a maximal matching $M$ of the edges in $G-E_{T}\subseteq G$ deterministically in $T_{\mathrm{MIS}}(\Delta\cdot n,2\Delta-2)$ rounds of the $\mathsf{LOCAL}$ model.
Proof 4.6.
A maximal matching of $G-E_{T}$ is simply an MIS on the line graph $L(G-E_{T})$.
4.3 Phase 2: Assign two exclusive edges to each cluster
Next, we rearrange the edges of $M$ in order to assign two exclusive edges to each cluster while maintaining the matching property. The two exclusive edges are needed in the subsequent steps to adjust the coloring of the intercluster edges such that the coloring can be completed within the cluster. This step is based on the HSO subroutine stated in Theorem B.8 and we describe in the following the specific instance that we use in our algorithm.
Algorithm 3 Flip$(\mathcal{C},M)$
1:Input: a $\Delta$-regular graph $G=(V,E)$, a clustering $\mathcal{C}$ and a matching $M\subseteq E$.
2:Output: a matching $M^{\prime}\subseteq E$ and an assignment $\psi:M^{\prime}\to\mathcal{C}^{\prime}$
3:
4:Each cluster $C^{-1}(r)\in\mathcal{C}^{\prime}$ finds an integer $k>0$ and a set of vertices $S\subseteq V(C^{-1}(r))$ of size at least $4\cdot\Delta^{3}$ such that $\mathrm{dist}(r,v)=k$ for all $v\in S$. $\triangleright$ Lemma 4.7
5:Construct auxiliary hypergraph $H=(\mathcal{C}^{\prime},M)$ with $\delta_{H}>\Delta\cdot r_{H}$. $\triangleright$ Lemma 4.10
6:Compute a $2$-HSO $\psi$ in $H$triangleright Lemma 4.12
7:Move matching edges closer to their cluster $\triangleright$ Lemma 4.14
8:return $M^{\prime},\psi$
The high level idea is that each cluster proposes to nearby edges in the existing maximal matching $M$. Each matching edge accepts exactly one proposal and we swap the edge towards the cluster that won it. This process can be modeled as an HSO instance where each matching edge in $M$ represents an hyperedge that contains as vertices all clusters that sent a proposal to that edge. Since a directed hyperedge $e$ is outgoing for exactly one of its vertices $v\in e$, we will say that $v$ won the proposal for $e$. However, the hypergraph sinkless orientation problem only guarantees at least one outgoing hyperedge for each vertex. Therefore, in order to achieve two exclusive hyperedges for each cluster, we split each cluster into two virtual nodes and let both participate in the HSO independently. In the following we formalize this idea and prove that this provides the desired properties as well as a deterministic runtime of $\mathcal{O}(\log n)$ rounds. First, we let each cluster $C^{-1}(r)\in\mathcal{C}$ that is not already edge-degree choosable select a subset of vertices $S\subseteq V(C^{-1}(r))$ to send proposals to a matching edges in their neighborhood. Let $\mathcal{C}^{\prime}$ denote the set of clusters, which are not already edge-degree choosable. Crucially, in order to satisfy the colorful condition from Lemma 3.4 we require that the vertices in $S$ are all leaves in the spanning tree $T(r)\subseteq C^{-1}(r)$ at the same distance from the root $r$.
Lemma 4.7.
For each cluster $C^{-1}(r)\in\mathcal{C}^{\prime}$, there exists a $k>0$ and a set of vertices $S\subseteq V(C^{-1}(r))$ of size at least $4\cdot\Delta^{3}$ such that
  • $\mathrm{dist}(r,v)=k$ for all $v\in S$ and
  • $v$ is a leaf in $T(r)$ for all $v\in S$.
Proof 4.8.
Since $\mathcal{C}$ is a $(4,8)$-clustering we have that $C^{-1}(r)$ necessarily contains all vertices in $N^{4}(r)$. Thus, according to Lemma 3.8 the corresponding BFS-tree contains at least $(\Delta-2)^{4}$ leaves. Therefore, by the pigeonhole principle there exists a layer $V_{k}$ for $k=1,\dots,8$ such that $V_{k}$ contains at least $(\Delta-2)^{4}/8>4\cdot\Delta^{3}$ leaves for $\Delta$ sufficiently large.
For each cluster we select an arbitrary set $S$ of at least $4\cdot\Delta^{3}$ leaves which are all at the same distance from the center. The subset $S$ is the subset of leaves that send requests. Each vertex in $v\in S$ sends exactly one proposal. If $v$ is incident to a matching edge $e\in M$, then it proposes to $e$. Otherwise, $v$ sends a proposal to an arbitrary matching edge that is one hop away from $e$. Such an edge must exist, due to the maximality of $M$.
Definition 4.9 (Auxiliary hypergraph).
We define the auxiliary hypergraph $H=(V_{H},E_{H})$:
  • For each cluster $C\in\mathcal{C}^{\prime}$ we add a vertex $v_{C}$ to $V_{H}$.
  • For each matching edge $e\in M$ we add a hyperedge $e_{H}$ to $E_{H}$ that contains all vertices $v_{C}$ for which there exists at least one vertex $v\in V(C)$ that proposed to $e$.
Lemma 4.10.
The minimum degree $\delta_{H}$ and the maximum rank $r_{H}$ of $H$ satisfy $\delta_{H}>\Delta\cdot r_{H}$.
Proof 4.11.
First we observe that a single matching edge $e=uv$ can get proposals only from $u,v$ and vertices that are exactly one hop away from either $u$ or $v$. Hence, there may be two proposals coming from $u$ and $v$ plus $2\cdot(\Delta-1)$ proposals coming from neighbors of $u$ or $v$. Thus, the maximum rank of $H$ is at most $r_{H}=2\Delta$. Next, each cluster contains at least $4\cdot\Delta^{3}$ vertices sending proposals. Since each vertex sends exactly one proposal and each edge receives at most $2\Delta$ proposals the vertices of each cluster must propose to at least $2\Delta^{2}$ edges in total. Hence, the minimum degree of $H$ can be lower bounded by $\delta=2\Delta^{2}$.
Lemma 4.12.
We can compute an orientation of $E_{H}$ such that each vertex in $V_{H}$ has at least two outgoing hyperedges in $\mathcal{O}(\log_{\Delta}n)$ rounds deterministically and $\mathcal{O}(\log_{\Delta}\log n)$ rounds randomized.
Proof 4.13.
To compute an orientation with two outgoing edges per cluster we simply split each vertex into two virtual nodes with half the degree each. Thus, the minimum degree of the resulting hypergraph $H^{\prime}$ is still at least $\Delta^{2}$. Therefore, we can solve HSO in $\mathcal{O}(\log_{\delta/r}n)=\mathcal{O}(\log_{\Delta}n)$ rounds deterministically and in $\mathcal{O}(\log_{\delta/r}\delta+\log_{\delta/r}\log n)=\mathcal{O}(\log_{\Delta}\log n)$ rounds randomized using the algorithms from Theorem B.8 and Theorem B.9.
The orientation of $H$ yields an assignment of matching edges to clusters such that each cluster receives exclusive access to two matching edges. However, the matching edges are not necessarily adjacent to their respective clusters yet. Therefore we modify the matching $M$ by moving each edge closer to the vertex which won the proposal for it.
Lemma 4.14.
There is a matching $M^{\prime}=\chi(M)$ such that for all $e\in M$ it holds that $\chi(e)$ is adjacent to the vertex that won the proposal for $e$.
Proof 4.15.
For each $e\in M$, let $v$ be the vertex (in the original graph $G$), which sent the winning proposal to $e$. If $v$ is incident to $e$, then we simply set $\chi(e)=e$. Otherwise, let $u$ be the endpoint of $e$ which is adjacent to $v$ and set $\chi(e)=uv$. With this procedure, we ensure that each cluster is adjacent to at least two edges in $M^{\prime}$ since each of the two partitions of the cluster receives at least one edge based on the HSO property. If a cluster is assigned more than two edges in $M^{\prime}$, it discards any edges beyond the required two. In order to show that the edge set $M^{\prime}$ is still a proper matching, we recall that every vertex $v\in S\subseteq V$ sends out exactly one proposal. Hence, there is no vertex that wins more than one edge in $M$. Assume that $v$ won the proposal for the edge $e=uw$. If $v$ is incident to $e$ there is no change from $M$ to $M^{\prime}$. Otherwise, let $u$ be adjacent to $v$. In this case, the edge $uw\in M$ is flipped to $vw\in M^{\prime}$. Hence, $v$ is the only vertex that gains an incident edge in $M^{\prime}$ compared to $M$. Since $e$ is one of the closest edges to $v$ in $M$ there cannot be an edge $e^{\prime}\in M$ that is already incident to $v$. Therefore every vertex that wins its only proposal is incident to exactly one edge in $M^{\prime}$. On the other hand, every vertex that does not win its proposal cannot gain any additional adjacent edges in $M^{\prime}$. Therefore $M^{\prime}$ is still a matching.
4.4 Phase 3: Switch colors in order to complete the coloring
Now that every cluster got access to two exclusive edges, which are at the same distance from the central vertex of the cluster, we show that we can modify these colors in order to satisfy the colorful condition from Lemma 3.4 for each cluster. To show that we can perform this procedure for all clusters in parallel, we argue that the colorful condition can still be fulfilled, even if the colors of all edges not adjacent to the two exclusive edges are modified adversarially.
Lemma 4.16 (Color switching).
Let $G=(V,E)$ be a $\Delta$-regular graph, $T=(V_{T},E_{T})\subseteq G$ a tree, $r\in V_{T}$ and $\varphi:E\setminus E_{T}\to[2\Delta-3]$ a partial edge coloring of $G$. Then, for any two independent edges $uv,u^{\prime}v^{\prime}\in E_{k}$, where $u$ and $u^{\prime}$ are leaves in $T$, there exists a coloring $\varphi^{\prime}:E\setminus E_{T}\to[2\Delta-2]$ such that $|\varphi^{\prime}(N_{E}(V_{k}))|\geq\Delta$ and $\varphi(e)=\varphi^{\prime}(e)$ for all $e\not\in\{uv,u^{\prime}v^{\prime}\}$.
(a) Before switching colors: $\varphi(N_{E}(V_{1}))=\{1,2\}$.
(b) After switching colors: $\varphi(N_{E}(V_{1}))=\{1,2,4\}$.
Figure 3: A simplified example for the color switching procedure, where $\Delta=3$. The edges in the BFS-tree $T_{r}$ are represented by solid lines, while edges outside of $T_{r}$ are dashed. The assigned matching edges $e_{1},e_{2}$ are marked orange.
Proof 4.17.
Recall that $V_{k}:=\{v\in V_{T}:\mathrm{dist}_{T}(v,r)=k\}$. Since $u$ and $u^{\prime}$ are leaves in $T$ they are incident to exactly $\Delta-1$ edges in $E\setminus E_{T}$. Therefore, $\varphi(N_{E}(V_{k}))$ already contains at least $\Delta-1$ distinct colors. If $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ we simply set $\varphi^{\prime}=\varphi$. Otherwise, if $\lvert\varphi(N_{E}(V_{k}))\rvert=\Delta-1$, we set $\varphi^{\prime}(uv)=2\Delta-2$. Since $\varphi(e)<2\Delta-2$ for all $e\in E_{k}$ we only need to argue that the color $\varphi(uv)$ still remains in $\varphi^{\prime}(E_{k})$. Consider the sets $N:=\{uw:uw\in E\setminus E_{T}\}$ and $N^{\prime}:=\{u^{\prime}w:u^{\prime}w\in E\setminus E_{T}\}$. Since $u,u^{\prime}$ are both leaves in $T$, both $N$ and $N^{\prime}$ have cardinality $\Delta-1$. Moreover, since $\lvert\varphi(E_{k})\rvert=\Delta-1$ and any two edges in $N$ (or $N^{\prime}$) are adjacent, we have that $\varphi(N)=\varphi(N^{\prime})$. Hence, there exists an edge $u^{\prime}v^{\prime\prime}$ such that $\varphi(uv)=\varphi(u^{\prime}v^{\prime\prime})=\varphi^{\prime}(u^{\prime}v^{\prime\prime})\in\varphi^{\prime}(E_{k})$. Thus, we have $\varphi^{\prime}(E_{k})=\Delta$. Note that this argument holds even if the colors of the edges not adjacent to $u$ or $v$ are modified adversarially.
Proof 4.18 (Proof of Theorem 1.1).
For $\Delta=\mathcal{O}(1)$ we can simply apply the $(3\Delta/2)$-edge coloring algorithm from Theorem B.12 to get a $(2\Delta-2)$-edge coloring in $\mathcal{O}(\log n)$ rounds. Hence, for the remainder of the proof we can assume $\Delta$ to be larger than a sufficiently large constant. Computing an MIS on the power graph $G^{8}$ takes $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds using the algorithm from Theorem B.10. Next, the clustering procedure can be done in a constant number of rounds, since the radius of each cluster is at most 8. For the $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$, as well as the maximal matching, we can again use the algorithm from Theorem B.10. Then, according to Section 4.3 we can rearrange the matching edges in a way that each cluster receives two exclusive edges to recolor. The rearrangement can be modeled as an instance of HSO with minimum degree $\frac{\Delta^{3}}{4}$ and maximum rank $\Delta^{2}$. The vertices of the virtual hypergraph are the clusters in $G$. Since each cluster has constant diameter, we can simulate the hypergraph with just a constant overhead factor. Thus, the deterministic HSO algorithm from Theorem B.8 runs in $\mathcal{O}(\log_{\Delta}n)$ rounds. Next we apply Lemma 4.16 to ensure that for each cluster there exists a $k$ such that $\lvert\varphi(N_{E}(V_{k}))\rvert\geq\Delta$ and by Lemma 4.16 we can extend $\varphi$ to a proper $(2\Delta-2)$-edge coloring of $G$.
We note that since the current state-of-the-art complexity for $(2\Delta-1)$-edge coloring and MIS coincide, this simple version of our algorithm already yields the $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ round algorithm claimed in Corollary 1.2. Our optimized version of the algorithm actually does not improve on the total runtime, but replaces the MIS and maximal matching subroutines by ruling sets and edge-ruling sets, respectively. Since $(2\Delta-1)$-edge coloring is in general a much easier problem than MIS, we expect that this optimization might lead to better runtimes in the future.
4.5 Removing MIS from our reduction
As it turns out, computing an MIS is not strictly necessary for our reduction. Indeed, using a $(2,k)$-ruling set just increases the diameter of our clusters. By tuning the parameter $k$ one finds that a $(2,\mathcal{O}(\log\Delta))$-ruling set is optimal for our purposes. This change only affects the runtime of some of our subroutines. For the second change, where we substitute the maximal matching by a $2$-edge ruling set, we actually need a slightly generalized version of how we set up our auxiliary hypergraph. We defer this modification to Section A.1 in the appendix. The changes to the simple version of Section 4.1 are also highlighted red in Algorithm 4.
Algorithm 4 Optimized $(2\Delta-2)$-edge coloring (high level overview)
1:Compute a $(2,\mathcal{O}(\log\Delta))$-ruling set $\mathcal{R}$ on $G^{8}$triangleright Corollary B.3
2:$\mathcal{C},E_{T}\leftarrow$ Cluster$(G,\mathcal{R})$triangleright Algorithm 2
3:Compute a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$triangleright Lemma 4.3
4:Compute a $2$-edge ruling set $E_{\mathcal{R}}$ of $T$triangleright Theorem B.7
5:Compute $2$-HSO $\psi:\mathcal{C}\to E_{\mathcal{R}}$ in $H=(\mathcal{C},E_{\mathcal{R}})$triangleright Theorem B.8
6:Rearrange the matching edges via $\chi:E_{\mathcal{R}}\to E_{\mathcal{R}}^{\prime}$triangleright Lemma 4.14
7:for each cluster $\mathfrak{C}\in\mathcal{C}$ in parallel do
8:     Obtain $\varphi^{\prime}$ from $\varphi$ by changing the colors of the edges in $\chi(\psi(\mathfrak{C}))$triangleright Lemma 4.16
9:end for
10:Extend $\varphi^{\prime}$ to a $(2\Delta-2)$-edge coloring $\varphi^{\star}$ of $G$triangleright Lemma 3.4
See 1.3
Proof 4.19.
The $(2,\mathcal{O}(\log\Delta))$-ruling set can be computed in $\mathcal{O}(\log\Delta+\log^{\ast}n)$ rounds according to Corollary B.3. As an immediate consequence, the diameter of our clusters now increases to $\mathcal{O}(\log\Delta)$. This implies that simulating the hypergraph now incurs a $\mathcal{O}(\log\Delta)$ factor overhead. Therefore the complexity of computing a HSO increases to $\mathcal{O}(\log_{\Delta}n\cdot\log\Delta)=\mathcal{O}(\log n)$ rounds. The $(2\Delta-3)$-edge coloring of $G-E_{T}$ remains unchanged. Further, we can also relax the requirements for our matching $M$ to $2$-edge ruling set, which can be computed in $\mathcal{O}(\log^{\ast}n)$ rounds. This requires us to be a little more careful in how we set up our hypergraph, which we defer to Section A.1 in the appendix. Therefore the total runtime of our reduction works out to be
$\mathcal{O}(\log\Delta+\log^{\ast}n)+\mathcal{O}(\log\Delta)+T_{2\Delta-1}(n)+\mathcal{O}(\log^{\ast}n)+\mathcal{O}(\log\Delta\cdot\log n)=\mathcal{O}(\log n)+T_{2\Delta-1}(n).$
5 Randomized $(2\Delta-2)$-edge coloring algorithm
In order to achieve a $\mathcal{O}(\log\log n)$-round randomized reduction, we only need to replace the ruling set algorithm by a faster randomized variant in Corollary B.5 and the deterministic HSO algorithm by its randomized counterpart in Theorem B.9. The changes to the optimized deterministic version are also highlighted red in Algorithm 5
Algorithm 5 Randomized $(2\Delta-2)$-edge coloring (high level overview)
1:Compute a $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set $\mathcal{R}$ on $G^{8}$triangleright Corollary B.5
2:$\mathcal{C},E_{T}\leftarrow$ Cluster$(G,\mathcal{R})$triangleright Algorithm 2
3:Compute a $(2\Delta-3)$-edge coloring $\varphi$ of $G-E_{T}$triangleright Corollary B.11
4:Compute a $2$-edge ruling set $E_{\mathcal{R}}$ of $T$triangleright Theorem B.7
5:Compute $2$-HSO $\psi:\mathcal{C}\to E_{\mathcal{R}}$ in $H=(\mathcal{C},E_{\mathcal{R}})$triangleright Theorem B.9
6:Rearrange the matching edges via $\chi:E_{\mathcal{R}}\to E_{\mathcal{R}}^{\prime}$triangleright Lemma 4.14
7:for each cluster $\mathfrak{C}\in\mathcal{C}$ in parallel do
8:     Change the colors of the edges in $\chi(\psi(\mathfrak{C}))$ if necessary $\triangleright$ Lemma 4.16
9:end for
10:Extend $\varphi$ to a $(2\Delta-2)$-edge coloring of $G$triangleright Lemma 3.4
See 1.5
Proof 5.1.
The $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set can be computed in $\mathcal{O}(\log\log n)$ rounds according to Corollary B.5. Hence, each cluster has diameter at most $\mathcal{O}(\log\log n)$ and can be computed in $\mathcal{O}(\log\log n)$ rounds as well. For the $2$-edge ruling set the deterministic algorithm from Theorem B.7 running in $\mathcal{O}(\log^{\ast}n)$ rounds is still fast enough for our purposes here. According to Theorem B.9 and Section 4.3 the hypergraph sinkless orientation can be computed in $\mathcal{O}(\log_{\Delta}\log n)$ rounds on the cluster graph, where we incur a $\mathcal{O}(\log\Delta)$-round simulation overhead. Finally, the extension to the edges inside each cluster can be computed in $\mathcal{O}(\log\log n)$ rounds. Hence, the total runtime works out to be
[…]
Similar to the deterministic case we could also get a $\mathcal{O}(\log_{\Delta}\log n)$-round reduction to randomized reduction via replacing $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set by an MIS. For randomized algorithms there is however a gap between the complexities for MIS and $(2\Delta-1)$-edge coloring. While $(2\Delta-1)$-edge coloring can be solved in $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds [20], applying the shattering framework of [18] to the deterministic MIS algorithm from [20] takes $\mathcal{O}(\log\Delta)+\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds. Hence, this alternative reduction does not yield any useful improvements at the moment.
6 Conclusion
Here we discuss some open questions and limits of our approach.
Why can we not use this approach for $\Delta$-vertex coloring?
The existing deterministic $\Delta$-coloring algorithms from the literature [37, 38, 22, 25] all follow a similar three-step approach: First they compute a ruling set $R$, then they color nodes outside $R$ in order of decreasing distance from $R$ and finally they color each vertex $v\in R$ via recoloring only a small-diameter subgraph around $v$. If we try to translate our approach into the vertex coloring setting we see that it is still possible to formulate an analogous condition for the extendability of partial vertex colorings to star graphs. However, clusters of larger diameter cannot be dealt with a similar approach. An intuitive reason for this is that the $\Delta$-vertex coloring is a much more restrictive problem. In order to find a color for the root $r$, it needs to have two neighbors $v_{1},v_{2}$ with the same color. In the literature this is often called having “slack” at $r$. If we now consider a $\Delta$-regular tree we unfortunately see that this conditions propagates through the layers of the tree, since it completely determines the colors of all vertices in the subtrees rooted at $v_{1}$ or $v_{2}$, up to permutations. One reason this approach works better for edge-coloring is the fact that one has just a bit more freedom in how one can place slack: If we want to color a tree from the leaves to the root, in the vertex coloring case, we need to place two nodes with the same color in layer $k+1$ to get a node with slack in layer $k$. In contrast, to obtain an edge with slack in layer $k$ we can use one edge from layer $k+1$ and the second one from layer $k$.
Appendix A Deferred proofs
A.1 Replacing the maximal matching by a $2$-edge ruling set
In this section we show that we can replace the maximal matching $M$ by a cheaper $2$-edge ruling set $E_{\mathcal{R}}$ of $G-E_{T}$. In this case, every leaf-vertex of a cluster sees at least one ruling set edge in its $3$-hop neighborhood. Here we need to be a little bit more careful in how we set up our hypergraph. Firstly, each vertex may only send one proposal to one of its closest ruling set edges. Secondly, each vertex may only propose to a ruling set edge $e\in E_{\mathcal{R}}$, if all the vertices on a shortest path towards this have proposed to $e$ as well. For an illustration we refer to Figure 4.
Figure 4: The ruling set edges are marked orange and dashed lines indicate the proposals.
For each cluster $C\in\mathcal{C}$ we select an arbitrary set $S(C)\subseteq V(C)$ of at least $4\cdot\Delta^{3}$ leaves which are all at the same distance from the center.
Definition A.1 (Proposal scheme).
Each cluster $C\in\mathcal{C}$ sends proposals to matching edges according to the following rules:
  • Each vertex $v\in S(C)$ proposes to exactly one matching edge $e\in E_{\mathcal{R}}$.
  • If $v$ proposes to $e$, there may not be an edge $e^{\prime}\in E_{\mathcal{R}}$ that is closer to $v$ than $e$.
  • If $v$ proposes to $e$, every vertex on the shortest path from $v$ to $e$ must also propose to $e$.
Definition A.2.
Let $e\in E_{\mathcal{R}}$ and $v\in V$ be the vertex which sent the winning proposal to $e$.
  • If $v$ is incident to $e$, then we simply add $e$ to $M^{\prime}$.
  • If $v$ is one hop away from $e$, let $u$ be the endpoint of $e$ which is adjacent to $v$ and add the edge $uv$ to $M^{\prime}$.
  • If $v$ is two hops away from $e$, let $u$ be the next vertex on a shortest path from $v$ to $e$ and add the edge $uv$ to $M^{\prime}$.
Now we prove an analogous result to Lemma 4.14.
Lemma A.3.
The edge set $M^{\prime}$ is a proper matching in $G-E_{T}$ and every edge $e\in M^{\prime}$ is adjacent to the vertex that won the proposal for $e$.
Proof A.4.
Recall that every vertex $v\in V$ sends out at most one proposal. Hence, there is no vertex that wins more than one edge in $E_{\mathcal{R}}$. Assume that $v$ won the proposal for the edge $e=uw\in E_{\mathcal{R}}$. If $v$ is incident to $e$ there is no change between $E_{\mathcal{R}}$ and $M^{\prime}$. If $v$ is one hop away from $e$, then $v$ is the only vertex that gains an incident edge in $M^{\prime}$ compared to $E_{\mathcal{R}}$. Since $e$ is one of the closest edges to $v$ in $E_{\mathcal{R}}$ there cannot be an edge $e^{\prime}\in E_{\mathcal{R}}$ that is already incident to $v$. In the final case that $v$ is two hops away from $e$, both $v$ and the next vertex $u$ on a shortest path from $v$ to $e$ gain an incident edge in $E_{\mathcal{R}}$. Note that every vertex is at most two hops away from its closest edge in $E_{\mathcal{R}}$, since $E_{\mathcal{R}}$ is a 2-edge ruling set. By our proposal scheme, $u$ must also have proposed to $v$. Therefore $u$ cannot have won its proposal and $u$ cannot be incident to an edge in $E_{\mathcal{R}}$. Hence, both $v$ and $u$ are incident to only one edge in $M^{\prime}$. Now assume that there is another vertex $v^{\prime}=v$ that adds the edge $v^{\prime}u$ to $M^{\prime}$. This means that $v^{\prime}$ must have proposed to the same edge as $u$. But $v$ already won the proposal for this edge, a contradiction! Therefore every vertex is incident to at most one edge in $M^{\prime}$.
Appendix B Subroutines
All subroutines from the existing literature used throughout our algorithms are stated here explicitly.
B.1 Ruling sets
In this subsection we collect and combine various ruling set algorithms from the distributed literature that we will use as subroutines for our edge coloring algorithm.
Theorem B.1 (ruling set algorithm [17, Theorem 8]).
There is a randomized distributed algorithm that computes a $(1,\mathcal{O}(\log\log\Delta))$-ruling set with induced degree at most $\mathcal{O}(\log^{5}n)$ in any $n$-vertex graph with maximum degree $\Delta$ in $\mathcal{O}(\log\log\Delta)$ rounds
Theorem B.2 (ruling set algorithm [41, Theorem 3]).
Let $G=(V,E)$ and $W\subseteq V$. Given a $d$-coloring of $G$, there is a deterministic distributed algorithm that computes a $(2,c)$-ruling set for $W$ in time $\mathcal{O}(c\cdot d^{1/c})$.
Corollary B.3.
There is a deterministic distributed algorithm that computes a $(2,\mathcal{O}(\log\Delta))$-ruling set in any graph $G$ in $\mathcal{O}(\log\Delta+\log^{\ast}n)$ rounds of the $\mathsf{LOCAL}$ model.
Proof B.4.
We start by computing a $\mathcal{O}(\Delta^{2})$-vertex coloring $\varphi$ of $G$ in $\mathcal{O}(\log^{\star}n)$ rounds using Linial’s algorithm [32, 33]. Then, we apply Theorem B.2 to $\varphi$ for $c=\mathcal{O}(\log\Delta)$ to get a $(2,\mathcal{O}(\log\Delta))$-ruling set in time $\mathcal{O}(c\cdot d^{1/c})=\mathcal{O}(\log\Delta\cdot\Delta^{2/\log\Delta})=\mathcal{O}(\log\Delta)$.
Corollary B.5.
There is a randomized distributed algorithm that computes a $(2,\mathcal{O}(\log\Delta)\cap\mathcal{O}(\log\log n))$-ruling set in any graph $G$ in $\mathcal{O}(\log\log n)$ rounds of the $\mathsf{LOCAL}$ model.
Proof B.6.
If $\Delta=\mathcal{O}(\log n)$, we can simply use the deterministic algorithm from Corollary B.3 to get an $(2,\mathcal{O}(\log\Delta))$-ruling set (which is also a $(2,\mathcal{O}(\log\log n))$-ruling set) in $\mathcal{O}(\log\Delta+\log^{\ast}n)=\mathcal{O}(\log\log n)$ rounds. Otherwise, for $\Delta=\Omega(\log n)$ we first apply Theorem B.1 to get a $(1,\mathcal{O}(\log\log\Delta))$-ruling set $S_{0}$ with induced degree at most $\mathcal{O}(\log^{5}n)$. Next we compute a $\mathcal{O}(\log^{10}n)$-coloring $\chi$ of $G[S_{0}]$ using Linial’s algorithm [32, 33]. Finally using $\chi$ as the input coloring and setting $c=\mathcal{O}(\log\log n)$, we get a $(2,\mathcal{O}(\log\log n))$ ruling set $S_{1}$ for $S_{0}$ in
$\mathcal{O}(c\cdot d^{1/c})=\mathcal{O}(\log\log n\cdot\log^{10/\log\log n}n)=\mathcal{O}(\log\log n).$
randomized time. Since every node $v\in V$ is at most $\mathcal{O}(\log\log\Delta)=\mathcal{O}(\log\log n)$ hops away from a vertex in $S_{0}$ and every vertex in $S_{0}$ is at most $\mathcal{O}(\log\log n)=\mathcal{O}(\log\Delta)$ hops away from a vertex in $S_{1}$, we get that $S_{1}$ is both a $(2,\mathcal{O}(\log\Delta))$-ruling set and a $(2,\mathcal{O}(\log\log n))$-ruling set for $V$.
Another subroutine that we will use in the optimized version of our algorithm computes a relaxation of a maximal matching.
Theorem B.7 ($2$-ruling edge set algorithm [31, Theorem 1.3]).
There is a deterministic distributed algorithm that computes a $2$-ruling edge set of $G$ in $\Theta(\log^{\star}n)$ rounds in the $\mathsf{CONGEST}$ model.
B.2 Hypergraph sinkless orientations
Theorem B.8 (deterministic HSO [8, Theorem 1.4]).
There is a deterministic distributed algorithm for computing an HSO in any $n$-vertex multihypergraph $H=(V,E)$ with maximum rank $r$ and minimum degree $\delta>r$ in $\mathcal{O}(\log_{\delta/r}n)$ rounds.
Theorem B.9 (randomized HSO [8, Theorem 1.5]).
There is a randomized distributed algorithm that w.h.p. computes an HSO on any hypergraph with maximum rank $r$ and minimum degree $\delta\geq 320r\log r$ with runtime $\mathcal{O}(\log_{\delta/r}\delta+\log_{\delta/r}\log n)$.
B.3 MIS & colorings
Theorem B.10 (deterministic MIS [20, Theorem 1.3]).
There is a deterministic distributed algorithm that in any $n$-vertex graph $G=(V,E)$ computes an MIS in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds of the $\mathsf{LOCAL}$ model. This complexity also applies for maximal matching, $(\mathrm{deg}+1)$-list vertex coloring, and $(2\cdot\mathrm{deg}-1)$-list edge coloring.
Corollary B.11 (randomized MIS [20, Theorem 1.3] & [10]).
There is a randomized distributed algorithm, in the $\mathsf{LOCAL}$ model, that computes a $(\Delta+1)$-vertex coloring in $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds. The same holds also for the $(2\Delta-1)$-edge coloring problem.
Theorem B.12 (deterministic $(3\Delta/2)$-edge coloring [8, Theorem 1.2]).
There is a deterministic $\mathcal{O}(\Delta^{2}\cdot\log n)$-round $\mathsf{LOCAL}$ algorithm that computes a $(3\Delta/2)$-edge coloring on any $n$-vertex graph with maximum degree $\Delta$.